# A Literature Review of Literature Reviews in Pattern Analysis and Machine Intelligence

### 20 Feb 2024 v1

@@ 1.1 Scope
In this study, we exclusively examine review articles. Grant’s comprehensive review [31] concluded that there are at least fourteen review types characterized by methods used, which highlights the fact that different reviews serve different purposes. For instance, a critical review strives to exhibit that the author has carried out extensive research on literature and evaluated it critically. A systematic review aims to systematically explore, assess and integrate research evidence, usually in adherence with review guidelines. However, these types of reviews are relatively rare in the field of PAMI. Given the predominance of literature reviews in this area, we entitle this paper ”A Literature Review of Literature Reviews.”
@@ 2.2.1 Data Source
RiPAMI(Ours) Complete Snapshots $\times$
@@
Reliable data sources for analyzing extensive reviews are fundamentally important. Based on the means of data acquisition and storage, existing scientific scholar data sources may be classified into two main categories: web-based and snapshot-based sources. Web-based source data refers to the meta-data that can be retrieved from the data provider in real-time with the use of the web crawler or the API (e.g. Semantic Scholar, arXiv, CrossRef). Such an approach would only consume a small amount of storage on the local machine but needs to query meta-data every single time. On the contrary, the offline snapshot consumes a larger amount of storage space but eliminates the need for frequent API queries. Since the snapshot is a mirror image of relevant papers before a specific time, its data remains unchanged over time compared to the API-based sources. As a result, snapshot ensures a consistent dataset in all of the experiments, avoiding issues of irreproducibility caused by changes in provided web retrieving service.
@@ 2.2.2 Database construction
Finally, a total of 2904 eligible papers with meta-data are retrieved and collected. To ensure reproducible experiments and prevent overburdening the server, we construct an SQL-based database dubbed RiPAMI (Reviews in Pattern Analysis and Machine Intelligence). This database stores information related to the paper such as title, abstract, date of publication, venue, citaion and reference details, etc.
@@ 2.2.3 Database Statistics
As aforementioned, The database consists of more than 2900 literature reviews from a variety of sources, publication years, and fields. To elucidate the characteristics of the RiPAMI database, we conduct a statistical analysis and plot the result in Fig. 3.
@@
Number of the References The number of references in a survey paper may influence its credibility and reliability. As shown in 3 (b), the distribution of literature review references in the RiPAMI database follows a log normal pattern. The average number of cited references is approximately 140, while the median number is 123.
@@ 4.1 Impact Indicators
TABLE II: Metrics for Evaluating Scholar Impact of Papers: “A” and “J” stand for article-level and journal-level. Filed normalized signifies that the metric can be utilized across fields.
@@ 4.1.2 Impact Evolution Index (IEI)
Most existing approaches treat the estimation of future citations as a sequence modeling task. Abrishami and Aliakbary propose employing the artificial neural network to predict long-term citations of a paper based on the number of its citations in the first few years after publication [1]. Zhao and Feng utilize graph structure representation and recurrent neural network modules to predict paper citation counts [120]. A more direct approach is to adopt polynomial fitting for scatter data, and calculate the sum of derivatives at each point. However, In practice, it has been observed that when dealing with data that lacks discernible distribution patterns, the polynomial fitting method tends to be sensitive to outliers, which can significantly compromise its robustness. Consequently, the numerical values obtained may not accurately reflect the underlying citation trend. In contrast to these series analysis-based methods, we propose a morphological theory-grounded Impact Evolution Index ($IEI$), which converts the citation trend into a clear and interpretable numerical value.
@@
$C^{\prime}_{i,n}(t)=n\sum_{i=0}^{n-1}\left(\binom{n-1}{i}(1-t)^{n-1-i}t^{i}\right)\left(P_{i+1}-P_{i}\right)$ (6)
Finally, the $IEI_{L_{l}}$ can be obtained by averaging the derivatives of $l=n+1$ points, as shown in Eq. (7). Note that the value of $l$ can be configured flexibly to meet actual demands. In general, the longer the period analyzed, the more stable the citation trend becomes. We usually prefer to analyze the most recent 6 months of citations when constructing a Bézier curve of degree 5. Furthermore, different months can contribute differently to the $IEI$. For instance, if we desire closer months to have a greater impact, we can achieve this by adjusting the weighting coefficients, $w_{i}$, of the derivatives at different points to calculate their weighted averages (See in Eq. (8)). In addition, the instantaneous trend could be regarded as the derivative of the last month in the sequence. It can be obtained by setting $w_{l}=1$ and the other weighting coefficients to 0. We denote the $IEI$ focused on the last month among the latest $l$ months (excluding the current month) as $IEI_{I_{l}}$ (See in Eq. (9)).
$IEI_{L_{l}}=(C^{\prime}_{0,l-1}(0)+\sum_{i=0}^{l-1-2}C^{\prime}_{i,l-1}(1))/(l-1)$ (7)
$IEI_{W_{l}}$ $=w_{0}C^{\prime}_{0,l-1}(0)+w_{1}C^{\prime}_{0,l-1}(1)$ (8)
$\quad+\ldots+w_{l}C^{\prime}_{l-1-2,l-1}(1)$
$IEI_{I_{l}}=C^{\prime}_{l-1-2,l-1}(1)$ (9)
In general, the longer the period being analyzed, the more stable the citation trend becomes. We usually prefer to analyze the most recent 6 months of citations when constructing a Bézier curve of degree 5 and utilize the curve to calculate $IEI_{L_{6}}$, $IEI_{W_{6}}$, and $IEI_{I_{6}}$.
@@ 4.2.2 Review Update Index (RUI)
The $RUI$ (Review Update Index) refers to the measure of the extent to which a literature review is required to be updated due to the iteration of technology, theory, etc. The index is related to both the literature itself and the research interests of the topic. Generally, a high update index suggests that a literature review is in need of an immediate update. Conversely, a lower update index implies that few advances have been made to the investigated field and the review is still up-to-date.
To evaluate the $RUI$, we may start with the coverage of references before and after publication. This coverage ratio can, to some extent, indicate the extent to which a review requires updating within its field. However, as mentioned earlier, accessing the coverage of a review is difficult. Fortunately, this problem is subtly avoided in calculating the ratio of relevant papers before and after publication. Assuming that the ratio of references containing the topic keyword in the title to all references is $R_{k}$, the total number of relevant articles can be estimated by dividing the number of articles containing those keywords retrieved from a search engine by $R_{k}$. Note that $R_{k}$ generally remains consistent before and after the publication, the CDR (Coverage Difference Ratio) can then be calculated in Eq. (13). The theoretical value range of CDR is greater than 0 to positive infinity. When the CDR of a review equals 1, it indicates that the current field has yielded new publications sufficient to constitute half of the literature referenced in the review.
@@
In addition, similar to the inevitable process of biological aging, literature reviews also undergo a gradual aging process throughout time. Such passage of time bestows upon literature reviews increasing aging progress, where the degree of aging can be conceptualized as a normalized value of the academic impact already achieved. To further explore the aging of reviews in the field of PAMI, we conducted a statistical analysis of the yearly number of newly received citations for reviews published between 2015-2017 in RiPAMI. In contrast to earlier findings, however, the distribution of received citations of reviews over time follows a t-distribution rather than a log-normal distribution of the regular paper, as previously reported in Ref. [66, 27, 71]. Due to the insufficient duration of published sample data, the observation of citation-time trend curves is incomplete. Therefore, we conducted a three-degree polynomial fitting on the limited 6-year citation trend data and transformed the positive segment of the fitted curve into a PDF. To obtain the corresponding cumulative distribution function (CDF), we employed the cumulative trapezoidal numerical integration method for an approximate estimation. Thus, the review aging degree is given by:
@@
[…]
@@ 4.3 Quantitative Evaluations of Literature Reviews
We have carefully selected numerous reviews within the field of PAMI to ensure a representative sample for our evaluation. By employing the proposed quantitative indicators, we conduct a thorough assessment of these scholarly surveys, aiming to gauge their quality and potential influence within the field. In Fig. 7, we present a set of scatter plots depicting the references’ quality of five randomly selected literature reviews published in two representative journals between 20015 and 2018. To prevent conflicts of interest, we use Journal A to represent a journal with an impact factor greater than 20. Journal B represents a journal with an impact factor below 5. The reference quality visualizations of journal A and journal B are presented in Fig. 7 (a), (b), respectively. The horizontal axis represents the reference age, and the vertical axis corresponds to the proposed $aTNCSI$. The color and size of the scatter points indicate the $IEI$ of the references. A positive $IEI$ value signifies a gradually increasing citation trend, resulting in warm-colored scatter points. The larger the size of the scatter point, the greater the value of $IEI$. Conversely, when the $IEI$ value is negative, the scatter points are cool-colored, and the size decreases as the value decreases. A closer examination of Fig. 7 reveals distinct distribution patterns among various journals depicted in their respective scatter plots. Journal with higher impact factors tends to exhibit clustered dots in the upper left corner, indicating a trend toward referencing more recent and influential sources. Conversely, journal with lower impact factors is prone to display more dispersed dots, with a notable distribution in the lower half of the graph. We argue that this discrepancy stems from authors’ tendency to cite recent papers in their literature reviews. Given that these relatively recent papers only had limited received citations, their citation count might not adequately reflect the potential value of the literature. Therefore, those seasoned researchers submitting reviews to top-tier journals often render more precise judgments about the prospective significance of a certain paper, as evidenced in the figure.
@@
More numerical results can be found in Tab. IV. As can be seen from the table, literature reviews [7, 18] of different topics that receive approximately the same number of citations exhibit notable discrepancies in their associated $aTNCSI$. This divergence primarily stems from the varying levels of research interest across their respective topics. Surveys [101] in some emerging fields reveal a significant increase in their $IEI$, as expected. The majority of leading journals demonstrate high $RQM$ values, which reflect the effectiveness of high-standard peer review. In addition, even if the topics covered by the reviews [119, 117] are similar and their publication dates are relatively close, the proposed $RUI$ still provide some indication of the extent to which they need to be updated. Note that, given the time span of less than six months, the calculation of $RUI$ for certain papers is unavailable (denoted with “-”).
@@ 5.1 Overview of AI-generated Literature Reviews
Though researchers are expecting to save significant time and effort in conducting comprehensive surveys by embracing AI-generated literature reviews, numerous concerns about ethical issues and fabrications in AI-generated academic content have been raised. A critical appraisal by Zybaczynska [124] highlights that current AI systems typically fall short in providing substantial, accurate information and critical discernment. Elali et al. [28] critically examines the profound challenges brought by the fabrication and falsification of AI-generated research, underscoring its significant impact on the scientific community. Furthermore, numerous academic publishers are cautious about content created by AI. For instance, journals such as “Science” and “Nature” prohibit the use of any text generated by ChatGPT or other automated tools in papers published in their issues. Nevertheless, we argue that the automated generation of literature reviews not intended for publication remains meaningful. It can assist researchers in staying rapidly abreast of the latest advancements in thriving fields.
@@ 5.2 Current State of AI-generated Literature Review
Despite previous discussions indicating that certain publishers refuse to accept content generated by LLMs, scholars have still successfully published AI-generated literature reviews in some of academic journals. Aydın et al. investigate how well can LLMs perform in the generation of the literature review. They employed well-known LLMs including ChatGPT and Google Bard to generate reviews on digital twin in healthcare [9] and meta-universes [8]. In their attempts, the Google Bard (or the ChatGPT) was first adopted to paraphrase the abstracts of papers within the last 3 years. Then, they designed several question prompts (such as “What is digital twin”) to query the LLMs and rearranged the response in a formatted style.

### 29 Feb 2024 v3

@@ 1.1 Scope
In this study, we exclusively examine review articles. Grant’s comprehensive review [31] concluded that there are at least fourteen review types characterized by methods used, which highlights the fact that different reviews serve different purposes. For instance, a critical review strives to exhibit that the author has carried out extensive research on literature and evaluated it critically. A systematic review aims to systematically explore, assess and integrate research evidence, usually in adherence with review guidelines. However, these types of reviews are relatively rare in the field of PAMI. Given the predominance of literature reviews in this area, we entitle this paper ”A Literature Review of Literature Reviews”.
@@ 2.2.1 Data Source
RiPAMI(Ours) Counts Only Snapshots $\times$
@@
Reliable data sources for analyzing extensive reviews are fundamentally important. Based on the means of data acquisition and storage, existing scientific scholar data sources may be classified into two main categories: web-based and snapshot-based sources. Web-based source data refers to the meta-data that can be retrieved from the data provider in real-time with the use of the web crawler or the API (e.g. Semantic Scholar, arXiv, CrossRef). Such an approach would only consume a small amount of storage on the local machine but needs to query meta-data every single time. On the contrary, the offline snapshot consumes a larger amount of storage space but eliminates the need for frequent API queries. Since the snapshot is a mirror image of relevant papers before a specific time, its data remains unchanged over time compared to the API-based sources. As a result, the snapshot ensures a consistent dataset in all of the experiments, avoiding issues of irreproducibility caused by changes in the provided web retrieving service.
@@ 2.2.2 Database construction
Finally, a total of 2904 eligible papers with meta-data are retrieved and collected. To ensure reproducible experiments and prevent overburdening the server, we construct an SQL-based database dubbed RiPAMI (Reviews in Pattern Analysis and Machine Intelligence). This database stores information related to the paper such as title, abstract, date of publication, venue, citation counts, and reference details, etc.
@@ 2.2.3 Database Statistics
The database consists of more than 2900 literature reviews from a variety of sources, publication years, and fields. To elucidate the characteristics of the RiPAMI database, we conduct a statistical analysis and plot the result in Fig. 3.
@@
Number of the References The number of references in a survey paper may influence its credibility and reliability. As shown in 3 (b), the distribution of literature review references in the RiPAMI database follows a log-normal pattern. The average number of cited references is approximately 140, while the median number is 123.
@@ 4.1 Impact Indicators
TABLE II: Metrics for Evaluating Scholar Impact of Papers: “A” and “J” stand for article-level and journal-level. Filed normalized signifies that the metric can be utilized across fields. Value normalized indicates that the range of the value is between 0 and 1.
@@ 4.1.2 Impact Evolution Index (IEI)
Most existing approaches treat the estimation of future citations as a sequence modeling task. Abrishami and Aliakbary propose employing the artificial neural network to predict long-term citations of a paper based on the number of its citations in the first few years after publication [1]. Zhao and Feng utilize graph structure representation and recurrent neural network modules to predict paper citation counts [120]. A more direct approach is to adopt polynomial fitting for scatter data, and calculate the sum of derivatives at each point. However, in practice, polynomial fitting methods tend to be sensitive to outliers when applied to data lacking discernible distribution patterns, potentially compromising their robustness significantly. Consequently, the numerical values obtained may not accurately reflect the underlying citation trend. In contrast to these series analysis-based methods, we propose a morphological theory-grounded Impact Evolution Index ($IEI$), which converts the citation trend into a clear and interpretable numerical value.
@@
$C^{\prime}(t)=n\cdot\sum_{i=0}^{n-1}B_{i,n-1}(t)\cdot\left(P_{i+1}-P_{i}\right)$ (6)
The tangent vector $C^{\prime}_{a}$ at the $a$-th point on the Bézier curve is further given by Eq. (7).
$C^{\prime}_{a}=$ $n\cdot\sum_{i=0}^{n-1}B_{i,n-1}(\frac{a}{n})\cdot\left(P_{i+1}-P_{i}\right),$ (7)
=\ = $(x_{a},y_{a}),\ a=0,1,\cdots,n$
where $x_{a}$ and $y_{a}$ are components of the vector, representing its magnitude along the x and y axes respectively.
Finally, the $IEI_{L_{l}}$ can be obtained by averaging the slope of $l=n+1$ distinct points on the curve, as shown in Eq. (8). Moreover, different months may contribute differently to the $IEI$. For instance, if we desire closer months to have a greater impact, we can achieve this by adjusting the weighting coefficients, $w_{a}$, of the slope at different points to calculate their weighted averages (See in Eq. (9)). In addition, the instantaneous trend could be regarded as the slope of the last month in the sequence. It can be obtained by setting $w_{n}=1$ and the other weighting coefficients to 0. We denote the $IEI$ focused on the last month among the latest $l$ months (excluding the current month) as $IEI_{I_{l}}$, as depicted in in Eq. (10).
$IEI_{L_{l}}=\sum_{a=0}^{n}\frac{(y_{a}/x_{a})}{n+1}$ (8)
$IEI_{W_{l}}=\sum_{a=0}^{n}\frac{w_{a}(y_{a}/x_{a})}{n+1}$ (9)
$IEI_{I_{l}}=C^{\prime}(1)=n\cdot(P_{n}-P_{n-1})$ (10)
Note that the value of $l$ can be configured flexibly to meet actual demands. In general, the longer the period being analyzed, the more stable the citation trend becomes. We usually prefer to analyze the most recent 6 months of citations when constructing a Bézier curve of degree 5 and calculate $IEI_{L_{6}}$, $IEI_{W_{6}}$, and $IEI_{I_{6}}$ based on the curve.
@@ 4.2.2 Review Update Index (RUI)
The $RUI$ refers to the measure of the extent to which a literature review is required to be updated due to the iteration of technology, theory, etc. The index is related to both the literature itself and the research interests of the topic. Generally, a high update index suggests that a literature review is in need of an immediate update. Conversely, a lower update index implies that few advances have been made to the investigated field and the review is still up-to-date.
To evaluate the $RUI$, we may start with the coverage of references before and after publication. This coverage ratio can, to some extent, indicate the extent to which a review requires updating within its field. However, as mentioned earlier, accessing the coverage of a review is difficult. Fortunately, this problem is subtly avoided in calculating the ratio of relevant papers before and after publication. Assuming that the ratio of references containing the topic keyword in the title to all references is $R_{k}$, the total number of relevant articles can be estimated by dividing the number of articles containing those keywords retrieved from a search engine by $R_{k}$. Note that $R_{k}$ generally remains consistent before and after the publication, the Coverage Difference Ratio ($CDR$) can then be calculated in Eq. (14). The theoretical value range of $CDR$ is greater than 0 to positive infinity. When the $CDR$ of a review equals 1, it indicates that the current field has yielded new publications sufficient to constitute half of the literature referenced in the review.
@@
In addition, similar to the inevitable process of biological aging, literature reviews also undergo a gradual aging process throughout time. Such passage of time bestows upon literature reviews increasing aging progress, where the degree of aging can be conceptualized as a normalized value of the academic impact already achieved. To further explore the aging of reviews in the field of PAMI, we conducted a statistical analysis of the yearly number of newly received citations for reviews published between 2015-2017 in RiPAMI. In contrast to earlier findings, however, the distribution of received citations of reviews over time follows a t-distribution rather than a log-normal distribution of the regular paper, as previously reported in Ref. [66, 27, 71]. Due to the insufficient duration of published sample data, the observation of citation-time trend curves is incomplete. Therefore, we conducted a three-degree polynomial fitting on the limited 6-year citation trend data and transformed the positive segment of the fitted curve into a PDF. To obtain the corresponding cumulative distribution function (CDF), we employed the cumulative trapezoidal numerical integration method for an approximate estimation. Thus, the Review Aging Degree (RAD) is given by:
@@
Figure 6: 3D visualization of the proposed quality indicators. Panels (a) and (b) respectively depict the value landscapes of $RQM$ and $RUI$, shaped by two independent variables.
@@ 4.3 Quantitative Evaluations of Literature Reviews
We have carefully selected numerous reviews within the field of PAMI to ensure a representative sample for our evaluation. By employing the proposed quantitative indicators, we conduct a thorough assessment of these scholarly surveys, aiming to gauge their quality and potential influence within the field. In Fig. 7, we present a set of scatter plots depicting the references’ quality of five randomly selected literature reviews published in two representative journals between 2015 and 2018. To prevent conflicts of interest, we use Journal A to represent a journal with an impact factor greater than 20. Journal B represents a journal with an impact factor below 5. The reference quality visualizations of journal A and journal B are presented in Fig. 7 (a), (b), respectively. The horizontal axis represents the reference age, and the vertical axis corresponds to the proposed $aTNCSI$. The color and size of the scatter points indicate the $IEI$ of the references. A positive $IEI$ value signifies a gradually increasing citation trend, resulting in warm-colored scatter points. The larger the size of the scatter point, the greater the value of $IEI$. Conversely, when the $IEI$ value is negative, the scatter points are cool-colored, and the size decreases as the value decreases. A closer examination of Fig. 7 reveals distinct distribution patterns among various journals depicted in their respective scatter plots. Journal with higher impact factors tends to exhibit clustered dots in the upper left corner, indicating a trend toward referencing more recent and influential sources. Conversely, a journal with lower impact factors is prone to display more dispersed dots, with a notable distribution in the lower half of the graph. We argue that this discrepancy stems from authors’ tendency to cite recent papers in their literature reviews. Given that these relatively recent papers only had limited received citations, their citation count might not adequately reflect the potential value of the literature. Therefore, those seasoned researchers submitting reviews to top-tier journals often render more precise judgments about the prospective significance of a certain paper, as evidenced in the figure.
@@
More numerical results can be found in Tab. IV. As can be seen from the table, literature reviews [7, 18] of different topics that receive approximately the same number of citations exhibit notable discrepancies in their associated $aTNCSI$. This divergence primarily stems from the varying levels of research interest across their respective topics. Surveys [101] in some emerging fields reveal a significant increase in their $IEI$, as expected. The majority of leading journals demonstrate high $RQM$ values, which reflect the effectiveness of high-standard peer review. In addition, even if the topics covered by the reviews [119, 117] are similar and their publication dates are relatively close, the proposed $RUI$ still provide some indication of the extent to which they need to be updated. Note that, given the time span of less than six months, the calculation of $IEI$ for certain papers is unavailable (denoted with “-”).
@@ 5.1 Overview of AI-generated Literature Reviews
Though researchers are expecting to save significant time and effort in conducting comprehensive surveys by embracing AI-generated literature reviews, numerous concerns about ethical issues and fabrications in AI-generated academic content have been raised. A critical appraisal by Zybaczynska [124] highlights that current AI systems typically fall short in providing substantial, accurate information and critical discernment. Elali et al. [28] critically examines the profound challenges brought by the fabrication and falsification of AI-generated research, underscoring its significant impact on the scientific community. Furthermore, numerous academic publishers are cautious about content created by AI. For instance, journals such as Nature” and Science” prohibit listing ChatGPT or other automated tools as authors on papers published in their issues. Nevertheless, we argue that the automated generation of literature reviews not intended for publication remains meaningful. It may assist researchers in staying rapidly abreast of the latest advancements in thriving fields.
@@ 5.2 Current State of AI-generated Literature Review
Despite previous discussions indicating that certain publishers refuse to accept non-human authors, scholars have still successfully published AI-generated literature reviews in some of academic journals. Aydın et al. investigate how well can LLMs perform in the generation of the literature review. They employed well-known LLMs including ChatGPT and Google Bard to generate reviews on digital twin in healthcare [9] and meta-universes [8]. In their attempts, the Google Bard (or the ChatGPT) was first adopted to paraphrase the abstracts of papers within the last 3 years. Then, they designed several question prompts (such as “What is digital twin”) to query the LLMs and rearranged the response in a formatted style.